# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v4

@@ 3.2 Dual-SVID Initialization
The resulting initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}=\mathrm{diag}(\mathbf{h}_{0})\mathbf{U}_{\mathrm{sign},0}\mathrm{diag}(\bm{\ell}_{0})\mathbf{V}_{\mathrm{sign},0}^{\top}\mathrm{diag}(\mathbf{g}_{0})$ preserves key structural information. The learnable parameters are initialized as follows: the latent factors $\mathbf{U}$ and $\mathbf{V}$ are initialized with the factors $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ obtained from SVD respectively, and the scales $\mathbf{h},\mathbf{g},\bm{\ell}$ are initialized to $\mathbf{h}_{0},\mathbf{g}_{0},\bm{\ell}_{0}$.
3.3 Residual Compensation
To further improve fidelity without increasing the primary path’s rank, we introduce Residual Compensation, a technique motivated by the theoretical insight that a two-stage error correction can be more effective than a single, larger approximation (Section˜A.1). Crucially, this method does not increase the model’s total parameter budget; instead, it strategically reallocates the fixed bit budget of a single, higher-rank approximation into two lower-rank paths: a primary and a residual path. The term ’residual’ primarily describes its role during initialization, where the auxiliary path is configured to model the error of the primary approximation. During QAT, both paths are optimized jointly to collectively represent the original weight. The advantage of this dual-path approach over a single path is empirically validated in our ablation studies (Section˜A.2.1).
@@ 3.3 Residual Compensation
This targeted error correction is often advantageous for maintaining performance at extreme compression levels, as demonstrated by the ablation studies presented in Section˜A.2.1. It is important to note that all learnable parameters from both the primary and residual paths are included when calculating the final BPW (see Section˜A.4 for calculation details).
Fig.˜3 visualizes the initial state of weight components ($\widehat{\mathbf{W}}_{\mathrm{pri},0}$, $\widehat{\mathbf{W}}_{\mathrm{res},0}$, and their sum $\widehat{\mathbf{W}}_{0}$) for a Llama2-7B query weight, derived via Dual-SVID across effective bits from 0.1 to 1.0, compared against the original weight $\mathbf{W}$ (far right). Even at a low 0.1 BPW, the primary approximation $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ (top row, leftmost) captures the dominant low-rank structure of $\mathbf{W}$, albeit with considerable simplification. As effective bits increase, $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ progressively refines details and more effectively suppresses disruptive approximation noise, allowing underlying weight patterns to emerge with greater clarity. A key observation is that $\widehat{\mathbf{W}}_{0}$ (bottom row) at a modest 0.3 BPW—benefiting from the initialized residual component $\widehat{\mathbf{W}}_{\mathrm{res},0}$ (middle row)—often presents a more faithful initial representation of $\mathbf{W}$ than $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ alone even at a higher 1.0 BPW. This underscores that Residual Compensation, even at the initialization stage, is vital for capturing complexities missed by the primary low-rank approximation, thereby providing a more faithful starting point for QAT.
@@ 4.1 Settings
We evaluate LittleBit on diverse LLM families including Llama [50], Llama2 [51], Llama3 [16], OPT [64], Phi-4 [1], and QwQ [49], spanning sizes from 1.3B to 32B parameters. The primary performance metric is perplexity (PPL), evaluated on the WikiText-2 [35] validation dataset. Additional results on the C4 [40] and PTB [33] datasets are provided in Section˜A.3. We also assess zero-shot accuracy on common reasoning benchmarks: WinoGrande [43], OpenBookQA (OBQA) [36], HellaSwag [59], BoolQ [5], ARC-Easy (ARC-e), ARC-Challenge (ARC-c) [6], and PIQA [4].
@@
The LittleBit model parameters, initialized using the Dual-SVID method (Section˜3.2), are optimized using QAT with knowledge distillation (KD) [19, 31, 65]. The original pre-trained full-precision model serves as the teacher ($\mathcal{T}$) for the LittleBit student model ($\mathcal{S}$). The QAT objective combines the standard output Kullback-Leibler (KL) divergence loss $\mathcal{L}_{\mathrm{out}}$ and an intermediate layer Mean Squared Error (MSE) loss $\mathcal{L}_{\mathrm{inter}}$ to match hidden representations, weighted by an empirically determined $\lambda=10$:
@@
Following the setup in [25], the training data combine WikiText-2 and selected partitions from C4. We use 2048 token sequence length, 5 epochs, Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$[28], and cosine LR decay with 2% warm-up (see details in  LABEL:appendix:hyper_parameters). For models employing Grouped-Query Attention (GQA) [2] (e.g., Llama3, Phi-4, QwQ), the latent ranks $r$ for key(K) and value (V) projection layers were specifically adjusted (see Section˜A.2.3) to maintain performance under extreme quantization. For backpropagation, SmoothSign (forward: $\mathrm{sign}(x)$; backward gradient: $\tanh(kx)$, $k=100$) is used, which generally performed better than Straight-Through Estimator (STE) [3] (see Section˜A.2.2). While the concept of a proxy gradient is not new, our contribution lies in the specific application and empirical validation of using the derivative of $\tanh(100x)$ as a stable gradient for QAT in the ultra-low-bit regime, which our ablation study confirms is effective (Section˜A.2.2).
@@
We benchmark LittleBit across a range of effective bits from approximately 1.0 down to 0.1 bits per weight (BPW). In the challenging sub-1-bit regime, we compare against STBLLM [10], a post-training quantization (PTQ) method employing N:M sparsity. For the $\sim$1.0 BPW setting, we compare against the 1-bit PTQ method BiLLM [20] and quantization-aware training (QAT) approaches (OneBit [57], BinaryMoS [25]). Results from standard low-bit methods like GPTQ [11] and OmniQuant [46] are included for broader context.
@@
4.2 Main Results
Superior Perplexity in Sub-1-bit Regime
Table˜1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [10], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
@@ 4.2 Main Results
A key strength of LittleBit is its exceptional stability and performance in the extreme sub-0.5 BPW range, specifically at 0.3 BPW and 0.1 BPW. In this territory, methods like STBLLM exhibit severe performance degradation (e.g., Llama2-7B PPL of 1.8e3 at 0.3 BPW). In contrast, LittleBit maintains strong PPL scores, such as 12.00 for Llama2-7B at 0.3 BPW and 15.92 at an unprecedented 0.1 BPW. When configured for an effective bits around 1.0 BPW by adjusting the latent rank $r$, LittleBit delivers perplexity comparable to specialized 1-bit QAT methods like OneBit [57] (e.g., Llama2-7B: LittleBit 9.08 vs. OneBit 8.36) and PTQ methods such as BiLLM [20]. While a method like BinaryMoS [25] may exhibit slightly lower PPL in some 1-bit scenarios (e.g., Llama2-7B: BinaryMoS 7.74), this can be attributed to its use of more complex mechanisms like dynamic scaling, which represent a different set of architectural and computational trade-offs. Analyzing the trade-off by adjusting latent rank to tune effective bits from 1.0 down to 0.1 BPW reveals that while performance degradation is minimal down to  0.55 BPW, a quantization cliff appears between 0.3-0.1 BPW, positioning the 0.3-0.55 BPW range as an optimal sweet spot for balancing compression and accuracy.
@@
LittleBit’s advantages are consistently observed across a diverse set of model architectures and sizes. This includes the Llama3 family, which is often considered challenging for quantization, and other contemporary models like Phi-4. For example, Llama3-8B quantized by LittleBit to 0.55 BPW yields a PPL of 18.47, significantly outperforming STBLLM’s 241.95 (at 0.55 BPW with 4:8 sparsity).The zero-shot evaluation results presented in Table˜2 further reinforce these perplexity-based findings. For instance, Llama2-7B compressed by LittleBit to 0.55 BPW achieves a mean accuracy of 47.26%, and even at an extreme 0.3 BPW, it maintains 45.20% mean accuracy. These results are notably competitive with, and in several cases surpass, those of STBLLM at similar or even higher effective bits (e.g., STBLLM Llama2-7B at 0.55 BPW scores 44.29%). This suggests a superior preservation of the models’ intrinsic reasoning capabilities even under extreme compression, highlighting LittleBit’s efficacy in creating highly compact yet powerful LLMs.
@@ 4.3 Reasoning Performance of Extremely Compressed LLMs
Figure 4: Zero-shot accuracy (%) on 7 common sense reasoning tasks for the Phi-4 14B model. Compares LittleBit-compressed Phi-4 at 0.55, 0.3, and 0.1 BPW against STBLLM.
The capability of LittleBit to achieve sub-1-bit compression while preserving significant model performance prompts an evaluation of their reasoning abilities, especially for large-scale models under extreme compression. We investigated the Phi-4 14B model, comparing its performance when compressed by LittleBit versus STBLLM across 7 zero-shot reasoning tasks. When compressed with LittleBit to 0.55 BPW, Phi-4 achieved an average accuracy of around 52.3%. While this performance is slightly lower than STBLLM at a similar compression level (0.55 BPW, 4:8 sparsity, achieving 54.2%), LittleBit demonstrates stronger graceful degradation. At a more aggressive 0.3 BPW, LittleBit-compressed Phi-4 maintained an average accuracy of 48.7%, significantly outperforming STBLLM at 0.3 BPW (2:8 sparsity), which scored 40.3%. Even at an extreme 0.1 BPW, Phi-4 compressed with LittleBit retained a notable average accuracy of 43.6%. These results highlight LittleBit’s ability to maintain more robust reasoning capabilities compared to STBLLM as compression becomes increasingly extreme, particularly in the sub-0.5 BPW regime for the Phi-4 model.
5 Analysis
Memory Footprint Reduction
[…]
@@ 5 Analysis
Table 4: Estimated KV cache memory reduction for Llama2-7B using LittleBit, with reduction factor $\sim d_{\mathrm{model}}/r$.
@@
[…]
@@
These results highlight LittleBit’s potential for efficient deployment with significant latency reduction. The observed gap between theoretical computational gains and practical latency stems from the fact that inference at small batch sizes is often dominated by memory access rather than raw computation. Our factorized method involves multiple memory-bound steps, and our kernel, while effective, is not as exhaustively optimized as industry-standard libraries. A detailed analysis including theoretical FLOPs, practical latency, and end-to-end decoding throughput is provided in LABEL:appendix:latency.
@@ 6 Conclusion
In this work, we proposed LittleBit, a novel method advancing LLM compression to the extreme sub-0.5 BPW regime, with viable performance at 0.1 BPW. Gains were achieved via SVD-inspired latent matrix factorization, binarization of these factors, and a multi-scale (row, column, latent) compensation mechanism. We also introduced Dual-SVID initialization for stable quantization-aware training (QAT) and integrated Residual Compensation for error mitigation. Empirical results show LittleBit significantly outperforms prior sub-1-bit techniques and maintains robust fidelity across LLMs up to 32B parameters. LittleBit thus establishes a new state-of-the-art in the size–performance trade-off, significantly expanding LLM deployment in resource-constrained environments. Building on this foundation, future work can enhance practical deployment through hardware co-design for edge devices like neural processing units (NPUs). Further gains in model fidelity can be explored by integrating advanced techniques like non-uniform bit allocation, applied either across different model layers or even within a single layer by dynamically distributing the bit budget between the primary and residual paths.
@@ Acknowledgements
Appendix A Appendix
A.1 Mathematical Analysis and Proofs
@@
Proposition 2 (Potential Advantage of Two‑Stage Quantization of SVD Components).
@@
A.2 Ablation Study
A.2.1 Residual Compensation

### v5

@@ 3.3 Residual Compensation
This targeted error correction is often advantageous for maintaining performance at extreme compression levels, as demonstrated by the ablation studies presented in Section˜B.1. It is important to note that all learnable parameters from both the primary and residual paths are included when calculating the final BPW (see Appendix˜D for calculation details).
Figure˜3 visualizes the initial state of weight components ($\widehat{\mathbf{W}}_{\mathrm{pri},0}$, $\widehat{\mathbf{W}}_{\mathrm{res},0}$, and their sum $\widehat{\mathbf{W}}_{0}$) for a Llama2-7B query weight, derived via Dual-SVID across effective bits from 0.1 to 1.0, compared against the original weight $\mathbf{W}$ (far right). Even at a low 0.1 BPW, the primary approximation $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ (top row, leftmost) captures the dominant low-rank structure of $\mathbf{W}$, albeit with considerable simplification. As effective bits increase, $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ progressively refines details and more effectively suppresses disruptive approximation noise, allowing underlying weight patterns to emerge with greater clarity. A key observation is that $\widehat{\mathbf{W}}_{0}$ (bottom row) at a modest 0.3 BPW—benefiting from the initialized residual component $\widehat{\mathbf{W}}_{\mathrm{res},0}$ (middle row)—often presents a more faithful initial representation of $\mathbf{W}$ than $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ alone even at a higher 1.0 BPW. This underscores that Residual Compensation, even at the initialization stage, is vital for capturing complexities missed by the primary low-rank approximation, thereby providing a more faithful starting point for QAT.
@@ 4.1 Settings
We evaluated LittleBit across diverse LLM families, including Llama [51], Llama2 [52], Llama3 [16], OPT [64], Phi-4 [1], and QwQ [50]. These models span parameter scales from 1.3B to 32B. Perplexity (PPL) on the WikiText-2 [36] validation dataset served as the primary performance metric. Appendix˜C provides additional results on the C4 [41] and PTB [33] datasets. Furthermore, we assessed zero-shot accuracy on common reasoning benchmarks: WinoGrande [44], OpenBookQA (OBQA) [37], HellaSwag [60], BoolQ [5], ARC-Easy (ARC-e), ARC-Challenge (ARC-c) [6], and PIQA [4].
@@
We optimized the LittleBit model parameters, initialized via the Dual-SVID method (Section˜3.2), using QAT with knowledge distillation (KD) [19, 31, 65]. The original pre-trained full-precision model functioned as the teacher ($\mathcal{T}$) for the LittleBit student model ($\mathcal{S}$). The QAT objective combines the standard output Kullback-Leibler (KL) divergence loss, $\mathcal{L}_{\mathrm{out}}$, and an intermediate layer mean squared error (MSE) loss, $\mathcal{L}_{\mathrm{inter}}$, to match hidden representations. We weighted these terms using an empirically determined coefficient $\lambda=10$:
@@
Adhering to the protocol in [25], the training data combined WikiText-2 with selected partitions from C4. The configuration included a sequence length of 2048 tokens, 5 epochs, the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$[28], and a cosine learning rate decay with 2% warm-up (see LABEL:appendix:hyper_parameters for details). For models employing Grouped-Query Attention (GQA) [2], such as Llama3, Phi-4, and QwQ, we specifically adjusted the latent ranks $r$ for key (K) and value (V) projection layers (see Section˜B.3) to maintain performance under extreme quantization. We employed SmoothSign (forward: $\mathrm{sign}(x)$; backward gradient: $\tanh(kx)$, $k=100$) for backpropagation, as it demonstrated superior performance compared to the Straight-Through Estimator (STE) [3] (see Section˜B.2). Although proxy gradients are established in the literature, this work validates the specific application of the $\tanh(100x)$ derivative as a stable gradient for QAT in the ultra-low-bit regime. Our ablation study confirms the effectiveness of this approach (Section˜B.2).
@@
We benchmarked LittleBit across effective bit rates ranging from approximately 1.0 down to 0.1 bits per weight (BPW). In the challenging sub-1-bit regime, we compared our method against STBLLM [10], a post-training quantization (PTQ) method that employs N:M sparsity. For settings approximating 1.0 BPW, we compared against the 1-bit PTQ method BiLLM [20] and quantization-aware training (QAT) approaches such as OneBit [58] and BinaryMoS [25]. We included results from standard low-bit methods, such as GPTQ [11] and OmniQuant [47], to provide broader context.
4.2 Main Results
Superior Perplexity in Sub-1-bit Regime
Table˜1 presents the perplexity (PPL) results of LittleBit compared to baseline methods. The proposed method demonstrates robust performance, particularly within the sub-1-bit regime. Relative to STBLLM [10], LittleBit achieves markedly superior PPL scores at $0.8$, $0.7$, and $0.55$ BPW. For instance, on Llama2-7B, LittleBit at $0.55$ BPW achieves a PPL of $10.47$, which represents a substantial improvement over the $30.67$ score reported for STBLLM. We observe comparable gains across other models and bit-widths. For Llama2-13B at $0.8$ BPW, LittleBit records a PPL of $8.52$, whereas STBLLM records $11.90$.
@@ 4.2 Main Results
(a) Phi-4 14B
(b) QwQ-32B
Figure 4: Zero-shot accuracy (%) on 7 common sense reasoning tasks. The comparison highlights the robustness of LittleBit (solid lines) against STBLLM (dashed lines) in the sub-0.5 BPW regime. While STBLLM degrades sharply, LittleBit preserves viable reasoning capabilities for both (a) Phi-4 and (b) QwQ models.
@@
A primary advantage of LittleBit lies in its exceptional stability and performance in the extreme sub-0.5 BPW range, specifically at $0.3$ BPW and $0.1$ BPW. In this regime, methods such as STBLLM exhibit severe performance degradation. For example, Llama2-7B yields a PPL of $1.8\times 10^{3}$ at $0.3$ BPW. Conversely, LittleBit maintains strong PPL scores, such as $12.00$ for Llama2-7B at $0.3$ BPW and $15.92$ at an unprecedented rate of $0.1$ BPW. When configured for effective bits approximating $1.0$ BPW by adjusting the latent rank $r$, LittleBit yields perplexity scores comparable to specialized 1-bit QAT methods. For example, on Llama2-7B, LittleBit achieves $9.08$ compared to $8.36$ for OneBit [58]. It also remains competitive with PTQ methods such as BiLLM [20]. Although BinaryMoS [25] may exhibit slightly lower PPL in certain 1-bit scenarios (e.g., Llama2-7B: BinaryMoS $7.74$), this performance difference is attributable to its use of more complex mechanisms, such as dynamic scaling, which represent a distinct set of architectural and computational trade-offs. Our analysis of the trade-off involved in adjusting the latent rank to tune effective bits from $1.0$ down to $0.1$ BPW indicates that performance degradation is minimal down to approximately $0.55$ BPW. However, a quantization cliff appears between $0.3$ and $0.1$ BPW. This observation positions the 0.30.30.550.55 BPW range as an optimal sweet spot for balancing compression and accuracy.
@@
We consistently observe the advantages of LittleBit across a diverse set of model architectures and sizes. This includes the Llama3 family, which is typically considered challenging for quantization, as well as other contemporary models such as Phi-4. For example, Llama3-8B quantized by LittleBit to $0.55$ BPW yields a PPL of $18.47$. This result substantially outperforms the STBLLM score of $241.95$ at $0.55$ BPW with 4:8 sparsity. The zero-shot evaluation results presented in Table˜2 corroborate these perplexity-based findings. For instance, Llama2-7B compressed by LittleBit to $0.55$ BPW achieves a mean accuracy of $47.26\%$, and the model maintains $45.20\%$ mean accuracy even at an extreme $0.3$ BPW. These results compare favorably with those of STBLLM at similar or even higher effective bits. STBLLM on Llama2-7B at $0.55$ BPW scores $44.29\%$. This outcome suggests superior preservation of the intrinsic reasoning capabilities of the models even under extreme compression. It highlights the efficacy of LittleBit in creating highly compact yet powerful LLMs.
@@ 4.3 Reasoning Performance of Extremely Compressed LLMs
The capability of LittleBit to achieve sub-1-bit compression while maintaining robust model fidelity prompts an evaluation of its reasoning abilities, particularly for large-scale models under extreme compression. We investigated the Phi-4 14B model and extended our analysis to the QwQ-32B model, comparing their performance against STBLLM across 7 zero-shot reasoning tasks.
When compressed with LittleBit to $0.55$ BPW, Phi-4 achieved an average accuracy of approximately $52.3\%$. Although this performance is marginally lower than STBLLM at a similar compression level ($54.2\%$ at $0.55$ BPW), LittleBit demonstrates substantially more graceful degradation. At the more aggressive $0.3$ BPW setting, the LittleBit-compressed Phi-4 maintained an average accuracy of $48.7\%$, which markedly outperforms STBLLM at $0.3$ BPW ($40.3\%$). Even at an extreme $0.1$ BPW, Phi-4 retained a notable accuracy of $43.6\%$. We observed a consistent trend with the QwQ-32B model, as illustrated in Figure˜4. While STBLLM exhibited severe performance collapse in the sub-0.5 BPW regime, LittleBit maintained a stable trajectory down to 0.1 BPW. These results highlight the ability of LittleBit to preserve robust reasoning capabilities across varying model scales as compression becomes increasingly extreme.
@@
Table 4: Estimated KV cache memory reduction for Llama2-7B using LittleBit, with a reduction factor of approximately $d_{\mathrm{model}}/r$.
@@
5 Analysis
Memory Footprint Reduction
LittleBit is designed for resource-constrained environments, such as on-device deployment, and achieves substantial reductions in model memory footprint. As summarized in Table˜3, quantizing a Llama2-7B model (originally $13.49$ GB in FP16) to an effective bit rate of $0.3$ BPW using LittleBit reduces its required storage to $0.79$ GB. This corresponds to a compression factor exceeding $17\times$. At the extreme setting of $0.1$ BPW, the footprint is further reduced to $0.63$ GB (over $21\times$ compression). We observe comparable reductions for larger models, such as Llama2-70B, where the memory footprint decreases from $138.04$ GB to under $2$ GB at $0.1$ BPW, achieving nearly $70\times$ compression. These substantial memory reductions significantly expand the range of devices capable of hosting large-scale language models, thereby enabling complex language tasks on hardware with limited VRAM or storage capacity.
@@ 5 Analysis
Figure 6: Kernel-level latency (bars, left axis) and speedup relative to FP16 (dashed line, right axis) on an NVIDIA A100 GPU for a Llama2-70B MLP layer ($8,192\times 28,672$). The benchmark compares the FP16 GEMM baseline with OneBit and the proposed LittleBit (LB) kernel across various BPW configurations ranging from 1.0 to 0.1.
@@
These findings suggest the proposed method enables efficient deployment with reduced latency. The discrepancy between theoretical computational gains and practical latency is likely attributable to memory access dominance during small-batch inference. The factorized approach involves multiple memory-bound operations, and the custom kernel has not yet reached the optimization level of industry-standard libraries. LABEL:appendix:latency provides a comprehensive analysis covering theoretical FLOPs, practical latency, and end-to-end decoding throughput.
@@ 6 Conclusion
This study proposed LittleBit, a method designed to advance LLM compression into the sub-0.5 BPW regime while maintaining viable performance at 0.1 BPW. Performance improvements were realized through SVD-inspired latent matrix factorization, factor binarization, and a multi-scale compensation mechanism encompassing row, column, and latent dimensions. Furthermore, Dual-SVID initialization was introduced to stabilize quantization-aware training (QAT), alongside Residual Compensation for error mitigation. Empirical results demonstrate that the proposed method markedly outperforms prior sub-1-bit techniques and preserves fidelity across LLMs scaling up to 32B parameters. Consequently, LittleBit suggests a favorable size–performance trade-off, thereby facilitating LLM deployment in resource-constrained environments. Future research may address practical deployment through hardware co-design for edge devices, such as neural processing units (NPUs). Additional improvements in model fidelity could be examined by integrating advanced techniques such as non-uniform bit allocation, applied either across model layers or within single layers via dynamic bit budget distribution between primary and residual paths.
@@ Acknowledgements
Appendix A Mathematical Analysis and Proofs
@@
Proposition 2 (Potential Advantage of Two-Stage Quantization of SVD Components).
@@
Appendix B Ablation Study
B.1 Residual Compensation